OpenStudy (anonymous):

3 5 - x + - = 4x find X ( slash marks are fractions 4 8

OpenStudy (austinl):

Subtract \(\frac{3}{4}x\) from both sides.

OpenStudy (anonymous):

fudge i got whats 5/8 minus 6/8???

OpenStudy (austinl):

You have this, \(\displaystyle \frac{6}{8}x+\frac{5}{8}=\frac{32}{8}x\) Subtract, 6/8 from 32/8. What do you get?

OpenStudy (anonymous):

26/8

OpenStudy (austinl):

Correct, so now we have, \(\displaystyle \frac{5}{8}=\frac{26}{8}x\) Now what would you do to solve for x now?

OpenStudy (anonymous):

its not letting me reply